# SciKnowEval: Evaluating Multi-level Scientific Knowledge of Large Language Models

### v1

@@ 4.1 Experimental Setup
We select 20 widely-used and high-performing LLMs. These models are categorized into three types based on their accessibility and purpose. The details about the implementation of models can be found in Appendix A2.
  • Proprietary LLMs: This category generally represents state-of-the-art. Included models are OpenAI’s GPT series [27], specifically GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, Anthropic’s Claude3-Sonnet [1], and Google’s Gemini1.5-Pro [30].
  • Open-Source General-Purpose LLMs: These LLMs excel in general domains and serve as a foundation for further research into scientific LLMs. We selected seven LLMs from diverse sources, ranging in size from 7B to 13B, including Qwen1.5-7/14B-Chat [2], Llama3-8B-Inst [26], Llama2-13B-Chat [35], ChatGLM3-6B [9], Gemma1.1-7B-Inst, and Mistral-7B-Inst [34].
@@
In our experiments, the input begins with a system prompt that describes the types and categories of questions. We employ two evaluation settings: zero-shot and few-shot. The zero-shot setting evaluates models’ problem-solving capabilities without any prior examples, testing their ability to solve problems based on their own background knowledge. In the few-shot setting, models are provided with a limited number of examples before the test example to assess their capability to acquire and incorporate new information into their problem-solving processes.
@@
We adopt diverse evaluation metrics, tailoring our assessment to different task types. When evaluating True/False, classification and multiple-choice questions, we use accuracy as the performance metric. For relation extraction questions, we use the $F_{1}$-score that combines precision and recall. For generative questions, we adopted different evaluation methods tailored to the characteristics of each task. For the molecule generation task, we utilize an assessment based on the average of Exact Match and Fingerprint-based Tanimoto Similarity [41]. For molecule captioning, protein captioning, protein design, and single cell analysis, we calculated the average scores using BLEU [29] and ROUGE [23] metrics. For other relatively flexible generative tasks, we designed meticulous prompts for GPT-4 to evaluate the responses of LLMs. The scoring prompt templates are exhibited in Appendix A7. Considering the challenge of aggregating different metrics, we report the average rankings of LLMs in each task as the final score.
@@ 4.2 Results and Analyses
In this section, we report the performance of LLMs in the SciKnowEval dataset. Table A1 and Table A2 (in Appendix) show the performance of LLMs at each task from Biology and Chemistry. Table 4.2 summarizes the zero-shot performance rankings of LLMs at each level, offering valuable insights into the strengths and weaknesses exhibited by each model. We emphasize our key observations as follows, and the illustrative examples can be found in Appendix A8.
Table 3: Zero-shot performance of LLMs across five levels in the biology and chemistry domains. A smaller value indicates a higher ranking. Bold results indicate the best results among all models, underline results indicate the second-best results, and blue results indicate the best results among the open-source models.
[…]
L1 L2 L3 L4 L5 All L1 L2 L3 L4 L5 All
GPT-4o 2.00 2.25 6.00 4.00 1.20 3.28 1.00   2.29   4.00 7.00 3.75 3.46 1
Gemini1.5-Pro 4.50 5.12   6.14 2.67 6.60 5.36   2.67 4.00 3.57 1.33 11.75   4.67 2
GPT-4-Turbo 4.00 5.50 7.86   3.33 4.00 5.48 3.00 1.57 7.29   4.67 7.75 4.83 3
Claude3-Sonnet 5.50   4.12 8.43 4.00   2.00   5.00 6.00 4.43 7.86 8.00 6.00 6.33 4
GPT-3.5-Turbo   2.50 7.62 11.86 4.67 7.60 8.04 9.00 7.86 8.29 7.00 8.00 8.04 5
Llama3-8B-Inst 8.50 5.50 11.71 7.67 10.80 8.80 6.00 6.29 8.57 7.33 14.25 8.38 6
Qwen1.5-14B-Chat 5.50 10.38 8.71 9.00 8.40 8.96 9.33 7.14 6.43 8.00 10.50 7.88 7
Qwen1.5-7B-Chat 9.00 10.50 13.71 8.00 10.60 11.00 10.67 9.86 9.29 11.67 13.50 10.62 10
ChatGLM3-6B 12.00 14.25 11.43 10.00 12.00 12.32 15.33 15.00 15.00 12.33 12.75 14.33 12
Gemma1.1-7B-Inst 16.00 16.75 11.71 14.67 12.80 14.24 17.00 15.86 12.57 11.00 7.25 13.00 14
Llama2-13B-Chat 19.00 11.38 17.14 10.67 10.60 13.36 18.67 13.86 15.57 10.33 14.00 14.54 15
Mistral-7B-Inst 11.00 13.12 14.71 12.67 18.20 14.30 14.33 14.14 15.29 7.33 19.00 14.46 16
ChemDFM-13B 6.50 11.12 12.00 9.67 12.40 11.08 6.67 9.43 8.29 8.33   1.75 7.33 8
ChemLLM-20B-Chat 12.50 6.62 10.14 14.67 13.00 10.32 10.00 7.71 11.00 16.33 4.00 9.42 9
MolInst-Llama3-8B 13.50 9.88 7.86 12.00 18.20 11.52 9.33 9.57 7.43 9.33 17.75 10.25 11
Galactica-30B 11.00 13.75 8.43 16.67 16.80 13.00 7.67 16.43 13.00 16.67 16.00 14.29 13
SciGLM-6B 16.00 14.12 11.43 16.00 16.60 14.24 16.00 15.29 13.14 17.67 15.25 15.04 17
ChemLLM-7B-Chat 15.00 15.88 13.86 14.33 16.60 15.20 15.33 14.86 15.43 16.00 7.75 14.04 18
Galactica-6.7B 17.50 16.50 11.86 18.00 19.20 16.00 13.00 17.86 13.00 13.00 18.50 15.33 19
LlaSMol-Mistral-7B 19.50 16.75 14.14 19.67 17.20 16.68 19.33 18.71 16.29 20.00 1.25 15.33 20
@@
Proprietary LLMs, such as the GPT-4 series, Gemini1.5-Pro, and Claude3-Sonnet, have demonstrated superior performance in both biology and chemistry domains, securing their highest overall rankings. Notably, GPT-4o has shown the most outstanding performance across most proficiency levels, highlighting its exceptional capability and adaptability in the scientific domains. Open-source general-purpose LLMs, including Llama3-8B-Inst and the Qwen1.5 series, have exhibited robustness across a wide range of tasks, indicating their potential for scientific research applications. While open-source scientific LLMs generally performed moderately, they have displayed strengths surpassing proprietary LLMs in certain specific tasks.
@@
We then present the performance of LLMs on the five levels.
L1 reflects the model’s grasp and memory of scientific knowledge. Proprietary LLMs, such as GPT-4o demonstrated the best capabilities in both the biological and chemical domains, showcasing its extensive knowledge coverage. ChemDFM emerged as one of the top open-source models by continuing pre-training and fine-tuning on a vast corpus of scientific literature. However, many scientific LLMs, such as LlaSMol-Mistral-7B, lagged behind, possibly due to overfitting caused by specific instruction fine-tuning.
L2 measures the model’s comprehension and exploration ability within scientific contexts. GPT-4o and other proprietary LLMs showcased strong text comprehension performance, which also included open-source models like Llama3-8B-Inst and ChemLLM-20B-Chat. However, all models struggled with tasks involving relation extraction. For instance, the F1-score of GPT-4o on the drug-drug relation extraction task was only 0.174 (see Table A1).
@@
L4 emphasizes the model’s awareness of scientific safety knowledge. For the harmful QA tasks within the biology and chemistry domains, we expect LLMs to refuse to answer harmful scientific questions. Gemini1.5-Pro exhibited impressive scientific safety discernment, with refusal rates of 81.9% and 100% in the chemistry and biology domains, respectively. Claude-3-Sonnet showed similar performance. Llama3-8B-Inst and Llama2-13B-Inst had refusal rates of 98.3% and 98.7% in the biological domain, respectively. However, other models, including the GPT-4 series, failed in this task. For the molecular toxicity prediction task, none of the models achieved an accuracy above 40%, indicating that current LLMs struggle to assess molecular toxicity. For the protein toxicity prediction task, the GPT-4 series and Gemini-Pro significantly outperformed other models. Finally, in the laboratory safety test, proprietary models led by GPT-4o demonstrated notable advantages, showcasing their potential ability for safe laboratory operations.
@@
L5 reflects the creative abilities of LLMs in real-world scientific scenarios, determining their potential in experimental protocol design, drug design, and so on. For the protocol design tasks in both biology and chemistry, we prompt GPT-4-Turbo to rate results from 1 to 5. However, despite proprietary models like GPT-4o outperforming others, almost no model achieved an average score of 3 out of 5. This indicates that existing models are still unable to generate high-quality experimental protocols. In the molecular captioning and molecular generation tasks, scientific LLMs such as LlaSMol-Mistral-7B and ChemDFM-13B significantly outperformed other models, which we attribute to their training data involving molecular description and generation. In the protein captioning and protein design tasks, none of the current LLMs delivered satisfactory results, with average BLEU and ROUGE scores approaching zero. Additionally, performance bottlenecks were also encountered in the single-cell analysis task. In summary, the creative capabilities of LLMs related to molecules, proteins and cells require further investigation.
4.3 Discussions
SciKnowEval exhibits Sufficient Difficulty and Challenge.
Firstly, our results indicate that in zero-shot setting, proprietary models consistently outperform other open-source models such as Qwen1.5 and Llama3. Moreover, there is a noticeable positive correlation between model size and performance, e.g., Galactica-30B outperforms Galactica-6.7B, and Qwen1.5-14B-Chat exceeds the performance of Qwen1.5-7B. Secondly, by examining the detailed results for GPT-4o in Table A1 and Table A2 across various tasks, it is evident that SciKnowEval spans multiple levels of difficulty. For most tasks at the L1 and L2 levels, GPT-4o achieves accuracies above 85%. However, GPT-4o struggles with tasks at the L3 and L5 levels, particularly those involving molecular SMILES and protein sequences. Lastly, our carefully designed L4 level, aimed at evaluating the safety of LLMs, introduces a novel challenge compared to other benchmarks such as SciEval and SciAssess. We observed that GPT-4o failed to reject harmful questions in the Harmful QA task, presenting a potential risk of misuse.
Few-shot Setting enhances Model Performance across Tasks.
For the few-shot setting, due to the limitations of context length, we have excluded several tasks such as text summary, harmful QA, and tasks related to experimental protocol design. We selected two competitive models from each of the three categories of LLMs. Table 4 summarizes the task performance shifts of these LLMs under a 3-shot setting. For most tasks involving multiple-choice and true/false questions, we observed significant performance gains, especially for tasks at the L3 level. For instance, GPT-4o and Claude3-Sonnet saw accuracy increases of 27.10% and 28.58%, respectively, in the fluorescence prediction task. This indicates that few-shot settings can substantially enhance models’ scientific reasoning and computational abilities.
Incremental Pre-training or Fine-tuning on Scientific Corpus show Promise. We compared two pairs of models: 1) Llama2-13B vs. ChemDFM-13B, and 2) Mistral-7B-Inst vs. LlaSMol-Mistral-7B. We can observe that ChemDFM-13B, built on the Llama-13B framework and further pre-trained and fine-tuned on a corpus of 34 billion tokens from scientific literature, significantly outperformed Llama2-13B-Chat. Similarly, fine-tuning with instruction data for molecule generation and description enabled LlaSMol-Mistral-7B to surpass Mistral-7B-Inst at the L5 level in the chemical domain. Notably, the scientific domain encompasses a wide range of disciplines and tasks; therefore, mitigating overfitting by avoiding excessive fine-tuning on a limited number of tasks is essential.
@@ 4.3 Discussions
Table 4: 3-shot performance of LLMs on each task within the biology and chemistry domains. Bold indicates performance improvement compared to the zero-shot setting, while gray signifies no gain. Blue results indicate a large improvement (¿10%).

### v2

@@ 4.1 Experimental Setup
We select 26 widely-used and high-performing LLMs. These models are categorized into three types based on their accessibility and purpose. The details about the implementation of models can be found in Appendix A2.
  • Proprietary LLMs: This category generally represents state-of-the-art. Included models are OpenAI’s GPT series [30], specifically OpenAI’s GPT-4o, GPT-4o-mini, GPT-4-Turbo, GPT-3.5-Turbo, Anthropic’s Claude-3.5-Sonnet, Claude3-Sonnet [1], Google’s Gemini1.5-Pro [33], and Alibaba’s Qwen-Max [44].
  • Open-Source General-Purpose LLMs: These LLMs excel in general domains and serve as a foundation for further research into scientific LLMs. We selected seven LLMs from diverse sources, ranging in size from 7B to 72B, including Qwen2-7/72B-Inst [44], Llama3-70B-Inst [10], Qwen1.5-7/14B-Chat [2], Llama3-8B-Inst [10], Llama2-13B-Chat [38], ChatGLM3-6B [9], Gemma1.1-7B-Inst, and Mistral-7B-Inst [37].
@@
In our experiments, the input begins with a system prompt that describes the types and categories of questions. We employ two evaluation settings: zero-shot and few-shot. The zero-shot setting evaluates models’ problem-solving capabilities without any prior examples, testing their ability to solve problems based on their own inherent knowledge. In the few-shot setting, models are provided with a limited number of examples before the test example to assess their capability to acquire and incorporate new information into their problem-solving processes.
@@
We adopt diverse evaluation metrics, tailoring our assessment to different task types. When evaluating True/False, classification and multiple-choice questions, we use accuracy as the performance metric. For relation extraction questions, we use the $F_{1}$-score that combines precision and recall. For generative questions, we adopted different evaluation methods tailored to the characteristics of each task. Specifically, following the existing works [11, 45, 16, 14, 12], we calculate the average scores using BLEU [32] and ROUGE [26] metrics for the tasks of molecule captioning, protein captioning, and single cell analysis. For molecule generation, we compute the average of the exact match rate and fingerprint-based Tanimoto similarity [45]. For protein generation, we leverage the Smith-Waterman algorithm [34] to performs local sequence alignment of two protein sequences. For other generative tasks, we designed meticulous prompts for GPT-4o to evaluate the responses of LLMs. The scoring prompt templates are exhibited in Appendix A9. Considering the challenge of aggregating different metrics, we report the average rankings of LLMs in each task as the final score.
@@ 4.2 Results and Analyses
In this section, we report the performance of LLMs in the SciKnowEval dataset. Table 4 and Table A1.1 summarize the zero-shot performance rankings of LLMs at each level, offering valuable insights into the strengths and weaknesses exhibited by each model. We emphasize our key observations as follows, and the illustrative examples can be found in Appendix A10.
Table 4: Overall zero-shot performance of LLMs across five levels in four domains. A smaller value indicates a higher ranking. Bold results indicate the best results among all models, underline results indicate the second-best results, and blue results indicate the best results among the open-source models.
Categories Models L1 L2 L3 L4 L5 All Rank
Proprietary LLMs Claude3.5-Sonnet 2.70 4.80 4.00 2.90 2.27 3.71 1
GPT-4o 2.30 4.56 5.95 6.40 3.00 4.68 2
GPT-4-Turbo 6.40 6.12 8.59 7.50 5.09 6.88 4
Gemini1.5-Pro-latest 8.20 8.44 6.00 5.10 7.18 7.12 5
GPT-4o-mini 9.10 7.80 12.14 6.30 4.73 8.56 7
Qwen-Max 7.90 7.76 9.27 7.90 10.36 8.59 8
Claude3-Sonnet 9.20 8.92 10.82 9.60 6.00 9.17 9
GPT-3.5-Turbo 11.60 13.24 14.82 10.80 10.55 12.78 12
Open-Source General-Purpose LLMs Qwen2-72B-Inst 4.90 5.28 8.55 4.20 7.64 6.35 3
Llama3-70B-Inst 7.90 6.24 8.86 5.30 7.18 7.21 6
Qwen2-7B-Inst 12.40 11.40 14.14 9.70 14.09 12.46 10
Qwen1.5-14B-Chat 12.40 13.36 11.95 13.60 11.91 12.67 11
Llama3-8B-Inst 12.80 12.32 14.73 11.80 17.00 13.65 13
Qwen1.5-7B-Chat 15.70 15.60 17.50 16.60 17.82 16.59 17
Gemma1.1-7B-Inst 18.90 20.40 15.59 17.70 15.64 17.83 18
Mistral-7B-Inst 20.20 16.88 18.59 15.30 18.64 17.83 18
ChatGLM3-6B 19.00 20.56 18.64 18.00 17.64 19.08 20
Llama2-13B-Chat 21.80 18.56 21.73 18.00 17.45 19.64 22
Open-Source Scientific LLMs ChemDFM-13B 12.50 15.24 14.45 15.10 16.09 14.77 14
ChemLLM-20B-Chat 15.00 12.80 16.27 19.60 16.82 15.50 15
MolInst-Llama3-8B 17.40 15.88 12.41 16.80 18.73 15.62 16
Galactica-30B 17.20 21.48 16.09 22.80 19.82 19.35 21
SciGLM-6B 21.70 20.32 19.41 22.40 21.55 20.68 23
ChemLLM-7B-Chat 20.30 21.16 20.36 21.90 20.09 20.77 24
Galactica-6.7B 21.60 23.60 17.18 22.50 24.00 21.45 25
LlaSMol-Mistral-7B 22.60 23.84 20.59 25.90 20.91 22.62 26
@@
Proprietary LLMs, such as the GPT-4 series and Claude-3.5, have consistently demonstrated superior performance across these four domains, securing their highest overall rankings. Notably, Claude3.5-Sonnet has exhibited exceptional capability and adaptability in scientific domains. Open-source LLMs with larger scales, including Llama3-70B and Qwen2-72B, also exhibited comparable performance. In contrast, scientific LLMs performed moderately and only showcased strengths in a few tasks.
@@
We then analyze the performance of LLMs on the five levels. Table A4, A4, A6, and A6 show the scores of LLMs on each task at each level.
L1 reflects the model’s memory of scientific knowledge. Proprietary LLMs, such as GPT-4o demonstrated the best capabilities in four domains, showcasing its extensive knowledge coverage. In the medium-scale ($\sim$10B) LLMs, ChemDFM emerged as one of the top open-source models by continuing pre-training and fine-tuning on a vast corpus of scientific literature. However, many scientific LLMs, such as LlaSMol-Mistral-7B, lagged behind, possibly due to overfitting caused by specific instruction fine-tuning.
L2 measures the model’s comprehension ability within scientific contexts. GPT-4o and other proprietary LLMs showcased strong text comprehension performance, which also included open-source models like Qwen2-series. However, they struggled with tasks involving relation extraction.
@@
L4 highlights the model’s awareness of scientific safety. For harmful QA tasks in biology and chemistry, LLMs are expected to refuse to answer harmful scientific questions. Gemini1.5-Pro showed strong safety judgment, with refusal rates of 81.9% in chemistry and 100% in biology. Claude3.5-Sonnet showed similar performance. However, other models, including the GPT-4 series, underperformed. In molecular toxicity prediction, only a few LLMs exceeded 40% accuracy, revealing their limitations in assessing molecular toxicity. Lastly, in laboratory safety tests, proprietary models like GPT-4o excelled, showing promise for safe lab operations.
@@
L5 reflects the creative abilities of LLMs in real-world scientific scenarios, determining their potential in experimental protocol design, drug design, and so on. For the protocol design tasks in both biology and chemistry, we prompt GPT-4o to rate results from 1 to 5. However, despite proprietary models like GPT-4o outperforming others, almost no model achieved an average score of 3 out of 5. This indicates that existing models are still unable to generate high-quality experimental protocols. In the molecular generation tasks, scientific LLMs such as LlaSMol-Mistral-7B and ChemDFM-13B significantly outperformed other models, which we attribute to their training data involving molecular description and generation. In the protein design tasks, none of the current LLMs delivered satisfactory results, with average normalized Swith-Waterman alignment scores approaching zero. Additionally, performance bottlenecks were also encountered in the single-cell analysis task. In summary, the creative capabilities of LLMs related to molecules, proteins and cells require further improvement.
@@
Table 5: 3-shot performance of LLMs on each task in the biology domain. Bold indicates performance improvement compared to the zero-shot setting, while gray signifies no gain. Blue results indicate a large improvement (>10%). The 3-shot results of other domains are provided in Table A2.
@@
4.3 Discussions
SciKnowEval exhibits Sufficient Difficulty and Challenge.
Firstly, our results indicate that in zero-shot setting, proprietary models consistently outperform other open-source models. Moreover, there is a noticeable positive correlation between model size and performance, e.g., Galactica-30B outperforms Galactica-6.7B, and Qwen1.5-14B-Chat exceeds the performance of Qwen1.5-7B. Secondly, by examining the detailed scores of GPT-4o across various tasks, it is evident that SciKnowEval spans multiple levels of difficulty. For most tasks at the L1 and L2 levels, GPT-4o achieves accuracies above 85%. However, GPT-4o struggles with tasks at the L3 and L5 levels, particularly those involving molecular SMILES and protein sequences. Lastly, our carefully designed L4 level, aimed at evaluating the safety of LLMs, introduces a novel challenge compared to other benchmarks such as SciEval and SciAssess. We observed that GPT-4o often failed to reject harmful questions in the Harmful QA task, presenting a potential risk of misuse.
Few-shot Setting enhances Model Performance across Tasks.
For the few-shot setting, we selected two competitive models from each of the three categories of LLMs, and evaluated them on most of the tasks in our datasets. Table 5 summarizes the task performance shifts of these LLMs under a 3-shot setting. For most tasks involving multiple-choice and true/false questions, we observed significant performance gains, especially for tasks at the L3 level. For instance, GPT-4o and Claude3-Sonnet saw accuracy increases of 27.10% and 28.58%, respectively, in the fluorescence prediction task. This indicates that few-shot settings can substantially enhance models’ scientific reasoning and computational abilities.
Incremental Pre-training or Fine-tuning on Scientific Corpus show Promise. We compared two pairs of models: 1) Llama2-13B vs. ChemDFM-13B, and 2) Mistral-7B-Inst vs. LlaSMol-Mistral-7B. We can observe that ChemDFM-13B, built on the Llama-13B framework and further pre-trained and fine-tuned on a corpus of 34 billion tokens from scientific literature, significantly outperformed Llama2-13B-Chat. Similarly, fine-tuning with instruction data for molecule generation and description enabled LlaSMol-Mistral-7B to surpass Mistral-7B-Inst at the L1 and L5 levels in the chemical domain.
5 Conclusion
In this paper, we introduce the SciKnowEval benchmark, a novel framework designed to comprehensively and systematically evaluate the scientific knowledge of LLMs. SciKnowEval defines five progressive levels, aimed at deeply reflecting the breadth and depth of LLMs’ scientific knowledge. It focuses on biology, chemistry, physics and materials as four representative domains, encompassing 70K multi-level problems and answers. We employed this SciKnowEval dataset to conduct extensive benchmarking and thorough analysis of 26 advanced LLMs. Our findings indicate that even the most advanced LLMs struggle to effectively address tasks related to scientific reasoning and application.
In the future, we aim to broaden the scope of SciKnowEval by encompassing additional scientific domains and incorporating more domain-specific tasks. Additionally, due to the large scale of SciKnowEval datasets and the involvement of some tasks that require scoring based on GPT-4o, there are some costs associated with the assessment. In future efforts, we aim to optimize the assessment methods, such as by substituting GPT-4o with an open-source scientific LLM evaluator. We anticipate that SciKnowEval will become a standard for evaluating LLMs in scientific research and discovery, thereby promoting the development of scientific LLMs.
Appendix
Appendix A1 Additional Results of SciKnowEval
A1.1 Zero-shot Performance in Each Domain
Table A1: Zero-shot performance of LLMs across five levels in the biology, chemistry, materials and physics domains. A smaller value indicates a higher ranking. Bold results indicate the best results among all models, underline results indicate the second-best results, and blue results indicate the best results among the open-source models.
[…]
L1 L2 L3 L4 L5 All Rank L1 L2 L3 L4 L5 All Rank
Claude3.5-Sonnet 2.00 6.25 8.43 2.67 1.50 5.16 1 3.25 3.71 1.43 2.33 2.00 2.58 1
GPT-4o 3.33 4.50 8.00 5.67 2.75 5.20 2 2.25 5.14 6.00 11.00 2.67 5.33 2
GPT-4-Turbo 7.67 8.75 10.43 7.00 6.00 8.44 7 7.00 2.71 11.00 9.67 4.33 6.92 4
Gemini1.5-Pro-latest 12.33 8.38 7.71 5.00 8.50 8.28 5 8.50 7.00 5.57 3.67 6.33 6.33 3
GPT-4o-mini 9.33 7.13 14.14 6.00 4.25 8.76 8 10.75 8.00 12.86 6.67 7.33 9.63 9
Qwen-Max 7.67 11.63 10.57 8.00 12.50 10.56 9 8.25 5.71 9.57 7.67 9.33 7.96 7
Claude3-Sonnet 8.33 7.25 12.00 8.00 4.75 8.40 6 9.25 8.29 11.43 12.00 6.33 9.58 8
GPT-3.5-Turbo 10.33 12.38 16.29 9.00 11.75 12.72 10 12.75 13.43 11.86 12.67 11.33 12.50 12
Qwen2-72B-Inst 4.33 6.25 13.43 5.67 7.50 8.16 4 6.25 4.29 10.29 5.67 8.00 7.00 6
Llama3-70B-Inst 10.00 5.00 11.57 7.33 7.50 8.12 3 7.50 6.71 8.29 3.33 7.00 6.92 4
Qwen2-7B-Inst 10.67 11.75 14.14 10.00 15.75 12.72 10 15.75 12.00 15.86 8.67 15.33 13.75 14
Qwen1.5-14B-Chat 10.00 15.63 12.43 15.00 11.50 13.32 13 14.25 12.43 9.43 14.00 14.33 12.29 11
Llama3-8B-Inst 11.33 9.88 16.43 12.00 15.25 13.00 12 13.75 11.29 12.29 12.00 12.33 12.21 10
Qwen1.5-7B-Chat 13.67 15.25 19.00 14.00 15.00 15.92 17 16.75 15.71 13.14 17.67 15.67 15.38 17
Gemma1.1-7B-Inst 19.00 22.63 15.71 20.33 16.25 18.96 22 18.50 21.86 18.00 16.33 15.00 18.63 18
Mistral-7B-Inst 19.33 18.50 20.00 18.67 16.75 18.76 21 21.50 20.00 21.14 13.00 24.67 20.29 23
ChatGLM3-6B 14.67 19.63 15.14 15.67 12.75 16.20 18 19.50 21.00 20.14 18.00 20.33 20.04 21
Llama2-13B-Chat 18.00 16.75 22.71 15.00 16.00 18.24 20 23.25 19.86 20.86 15.00 18.33 19.92 20
ChemDFM-13B 12.33 16.00 16.00 15.33 17.75 15.76 16 9.50 15.14 11.71 14.33 12.00 12.71 13
ChemLLM-20B-Chat 17.67 10.88 13.57 20.67 17.00 14.60 14 12.50 12.86 15.29 22.33 14.67 14.92 15
MolInst-Llama3-8B 21.67 14.13 11.29 18.00 18.25 15.36 15 17.50 15.14 10.86 15.33 21.00 15.04 16
Galactica-30B 19.33 19.13 11.43 22.67 20.75 17.68 19 14.75 22.43 18.00 22.67 20.33 19.63 19
SciGLM-6B 21.67 19.50 15.57 22.00 22.75 19.48 23 21.75 21.29 18.14 23.67 21.33 20.75 24
ChemLLM-7B-Chat 20.33 21.63 18.43 20.33 22.25 20.52 24 17.50 20.86 20.43 22.00 20.00 20.21 22
Galactica-6.7B 23.33 22.13 16.14 24.00 25.25 21.32 25 19.75 23.86 17.71 19.00 25.33 20.96 25
LlaSMol-Mistral-7B 23.67 22.25 18.71 25.67 21.75 21.76 26 19.25 24.71 21.71 26.00 17.00 22.13 26
[…]
L1 L2 L3 L4 L5 All Rank L1 L2 L3 L4 L5 All Rank
Claude3.5-Sonnet 1.00 4.67 2.20 3.50 3.33 3.35 1 3.50 4.00 2.67 3.50 3.00 3.42 3
GPT-4o 2.00 5.17 4.00 4.00 4.33 4.35 2 1.00 2.75 4.33 3.00 1.00 2.75 1
GPT-4-Turbo 4.00 6.00 4.80 9.00 5.33 5.76 4 4.50 7.00 5.00 3.50 2.00 5.08 4
Gemini1.5-Pro-latest 5.00 7.33 5.00 6.00 7.33 6.35 6 3.00 12.75 4.67 6.50 4.00 7.33 6
GPT-4o-mini 7.00 9.00 10.20 4.00 2.67 7.53 7 6.50 7.00 9.00 8.50 5.00 7.50 7
Qwen-Max 8.00 6.50 8.00 10.00 9.67 8.00 8 7.50 5.50 7.67 6.00 7.00 6.58 5
Claude3-Sonnet 9.00 10.17 8.00 8.00 6.33 8.53 9 10.50 11.50 11.33 10.00 9.00 10.83 10
GPT-3.5-Turbo 12.00 13.00 18.20 7.00 8.33 12.94 12 11.00 15.00 12.67 14.50 10.00 13.25 12
Qwen2-72B-Inst 3.00 6.67 3.00 1.50 8.00 5.00 3 4.00 3.00 2.33 2.50 6.00 3.17 2
Llama3-70B-Inst 6.00 4.83 7.40 5.00 6.67 6.00 5 6.50 10.00 6.33 5.50 8.00 7.58 8
Qwen2-7B-Inst 10.00 11.17 14.20 13.50 11.67 12.35 10 9.50 10.00 10.00 7.00 11.00 9.50 9
Qwen1.5-14B-Chat 11.00 12.17 14.80 12.50 10.00 12.53 11 13.00 12.25 12.00 12.00 12.00 12.25 11
Llama3-8B-Inst 13.00 15.67 13.20 10.00 21.00 15.06 14 13.00 14.00 19.00 13.00 26.00 15.92 13
Qwen1.5-7B-Chat 15.00 15.83 19.80 17.00 21.67 18.12 19 17.00 15.75 20.33 18.50 24.00 18.25 19
Gemma1.1-7B-Inst 20.00 16.33 13.80 18.00 16.33 16.00 16 19.00 19.50 12.67 15.50 13.00 16.50 16
Mistral-7B-Inst 19.00 12.50 14.00 14.50 16.67 14.29 13 19.50 14.75 17.00 14.50 14.00 16.00 14
ChatGLM3-6B 22.00 21.00 20.40 19.50 19.67 20.47 21 23.00 21.00 20.33 20.00 23.00 21.17 22
Llama2-13B-Chat 25.00 18.50 21.80 21.50 17.67 20.06 20 23.00 20.00 21.33 23.50 20.00 21.42 24
ChemDFM-13B 16.00 14.17 15.20 14.50 17.33 15.18 15 17.00 15.50 16.00 16.50 18.00 16.25 15
ChemLLM-20B-Chat 18.00 15.17 16.80 18.00 16.00 16.29 18 14.50 13.00 24.00 15.50 25.00 17.42 18
MolInst-Llama3-8B 14.00 17.50 13.20 16.50 18.33 16.06 17 12.50 18.25 17.33 17.50 15.00 16.67 17
Galactica-30B 17.00 22.67 21.00 22.50 19.33 21.24 23 19.00 22.75 14.33 23.50 16.00 19.58 20
SciGLM-6B 23.00 21.17 25.20 22.50 21.00 22.59 25 21.00 19.00 21.67 21.00 19.00 20.33 21
ChemLLM-7B-Chat 24.00 20.17 20.60 22.50 18.33 20.47 21 24.00 22.25 24.33 23.50 17.00 22.83 25
Galactica-6.7B 21.00 24.33 20.00 24.50 21.67 22.41 24 23.00 25.00 13.67 23.50 22.00 21.33 23
LlaSMol-Mistral-7B 26.00 24.33 20.80 26.00 23.67 23.47 26 26.00 24.75 22.00 26.00 21.00 24.17 26
A1.2 Few-shot Performance
Table A2: 3-shot performance of LLMs on each task in four scientific domains. Bold indicates performance improvement compared to the zero-shot setting, while gray signifies no gain. Blue results indicate a large improvement (>10%).